# Creating Multi-Level Skill Hierarchies in Reinforcement Learning

### v1

@@ 6 Discussion and Future Work
Appendix A Sensitivity to the Resolution Parameter
Here we explore how changing the resolution parameter $\rho$ impacts the partitions produced by the Louvain algorithm and the performance of agents with the Louvain skill hierarchy.
The resolution parameter $\rho$ controls the relative importance placed on the two terms that make up the definition of modularity: (1) $e_{ii}$, the proportion of total edge weight in the graph that connects two nodes that are in the same cluster $c_{i}$, and (2) $a_{i}$, the proportion of total edge weight in the graph where at least one end of the edge is a node in cluster $c_{i}$. A lower value of $\rho$ favours the $e_{ii}$ term, rewarding large clusters with many intra-cluster edges. This leads the Louvain algorithm to run for more iterations, producing more partitions and resulting in a deeper skill hierarchy. Conversely, a higher value of $\rho$ favours the $a_{i}$ term, penalising inter-cluster edges and overall cluster size more harshly. This leads the Louvain algorithm to run for fewer iterations, producing fewer partitions and resulting in a shallower skill hierarchy.
Figure 1: How the output of the Louvain algorithm changes with the resolution parameter in Rooms and Towers of Hanoi.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: How the output of the Louvain algorithm changes with the resolution parameter in Rooms and Towers of Hanoi.
Figure 2: How agent performance changes when using Louvain skills created using different values of the resolution parameter $\rho$.
Figure 2 shows how changing $\rho$ impacts the partitions produced in Rooms and Towers of Hanoi. At higher $\rho$ values, fewer partitions are produced, each containing many small clusters. As $\rho$ decreases, more partitions are produced because the Louvain algorithm performs additional iterations and merges the clusters found in earlier partitions to produce larger clusters. For example, in Towers of Hanoi, at $\rho=10$, only a single partition containing many small clusters is produced. At $\rho=0.1$, three partitions are produced, each containing progressively fewer, larger clusters. As $\rho$ decreases, the clusters identified at a given level of the hierarchy generally remain stable: the lowest-level partition found using a higher value of $\rho$ will typically be the same as the lowest-level partition found using a lower value. In other words, decreasing $\rho$ may add levels to a hierarchy but it generally does not impact the existing levels of the hierarchy.
When using extremely high values of $\rho$, the Louvain algorithm terminates without producing any partitions. In contrast, when using extremely low values of $\rho$, the Louvain algorithm runs until a partition is produced where all nodes are merged into a single cluster.
@@ Appendix A Sensitivity to the Resolution Parameter
Figure 2 shows the performance of agents in Rooms and Towers of Hanoi using Louvain skills created by using different values of $\rho$. We found that Louvain skills created using lower values of $\rho$ consistently outperformed those created by using higher values, and that very high values ($\rho\geq 10$) generally led to performance similar to that obtained by using primitive actions only. This makes sense because lower $\rho$ values lead to deeper hierarchies that contain skills for navigating the environment over varying timescales. In contrast, higher $\rho$ values result in shallower skill hierarchies that contain few—or, in the extreme, no—levels of skills above primitive actions. While there are better and worse values of $\rho$, it may be argued that there is no “bad” choice; lowering $\rho$ will result in deeper hierarchies, but existing levels of the hierarchy will remain intact, and all levels will be made up of skills that are useful and intuitively appealing.
Appendix B Louvain Algorithm Pseudocode
@@
Algorithm 1 shows pseudocode for the Louvain algorithm. It takes a graph $G_{0}=(V_{0},E_{0})$ as input and outputs a set of partitions of that graph into clusters, with one partition being produced for each iteration before termination. This series of partitions forms the basis of our characterisation of a useful skill hierarchy for reinforcement learning.
@@
1 parameters: resolution parameter $\rho\in\mathbb{R}^{+}$ input: $G_{0}=(V_{0},E_{0})$ // e.g., the state-transition graph (STG) of an MDP $i\leftarrow 0$ repeat
@@ 8             $i\leftarrow i+1$
Algorithm 1 The Louvain Algorithm for Hierarchical Graph Clustering
Appendix C Incremental Method Pseudocode
Algorithm 2 shows pseudocode for a high-level algorithm where an agent, starting with only primitive actions and no information about its environment, incrementally builds a state-transition graph and creates Louvain skills from it.
@@ Appendix C Incremental Method Pseudocode
$G\leftarrow(V,E)$   // initialise empty state-transition graph (STG)
@@ 20            $t\leftarrow t+1$
Algorithm 2 Incrementally Discovering and Using Louvain Skills
Algorithm 3 shows an incremental version of the Louvain algorithm that integrates new nodes into an existing Louvain cluster hierarchy.
parameters: resolution parameter $\rho\in\mathbb{R}^{+}$ input: $G_{0}=(V_{0},E_{0})$   // e.g., the state-transition graph (STG) of an MDP
@@ 8             $i\leftarrow i+1$
Algorithm 3 Incremental Louvain Algorithm for Hierarchical Graph Clustering
Appendix D Environments
Gridworlds. In Rooms, Grid, Maze (Ramesh et al., 2019), and Office, the agent learns to navigate from a starting grid-square to a goal grid-square. Four primitive actions are available in each state: north, south, east, and west. Moving into a wall causes the agent to remain in the same state. The agent receives a reward of $-0.001$ per action taken, and an additional $+1.0$ for reaching the goal state.
Multi-Floor Office. Multi-Floor Office is an extension of the Office environment to multiple floors, with each floor having a central elevator grid-square that allows the agent to move between adjacent floors. When standing on an elevator grid-square, the agent has access to two additional primitive actions: up and down. The dynamics of Multi-Floor Office are otherwise identical to Office.
Taxi. The taxi agent transports passengers between four taxi-ranks (R, G, B, and Y) in a gridworld Dietterich (2000). The agent starts in a random grid-square, with a passenger waiting at one taxi-rank to be transported to a different taxi-rank. Six primitive actions are available in each state: north, south, east, west, pick-up-passenger, and put-down-passenger. Moving into a wall or inappropriately trying to pick up or put down a passenger cause the agent to remain in the same state. The agent receives a reward of $-0.001$ for each action, and $+1.0$ for delivering a passenger to their destination taxi-rank.
Towers of Hanoi. An agent starts with four different-sized discs stacked on the leftmost of three pegs, and its goal is to move them to all be stacked on the rightmost peg. In each state, the agent has access to primitive actions that move the top disc from one pole to any other pole, with the constraint that a larger disc cannot be placed on top of a smaller disc. The agent receives a reward of $-0.001$ for action action, with an additional reward of $+1.0$ for reaching the goal state.
Appendix E Full Cluster Hierarchies Produced in Sample Domains
Here we show the Louvain cluster hierarchies in Grid and Maze environments. We also show the first level of the hierarchy in Office, which is omitted in the main paper due to space limitations.
Level 4
Level 3
Level 2
Level 1
Figure 3: Partitions produced in Grid.
Level 5
Level 4
Level 3
Level 2
Level 1
Figure 4: Partitions produced in Maze.
Level 1
Figure 5: Lowest-level partition produced in Office.
Appendix F Methodology
How we produced options in each environment. For Louvain options, we started by applying the Louvain algorithm ($\rho=0.05$) to the state-transition graph. We discarded any lower-level partition that had a mean number of nodes of less than $4$. For each cluster $c_{i}$ in the remaining partitions, we defined options for navigating to each neighbouring cluster $c_{j}$.
To produce skills using the Label Propagation Farahani and Mozayani (2019), Edge Betweenness Davoodabadi and Beigy (2011), and Xu et al. (2018) methods, we applied the appropriate graph partitioning method to the state-transition graph and defined skills for moving between neighbouring clusters in the resulting partition.
We produced Eigenoptions Machado et al. (2017) by computing the normalised Laplacian of the state-transition graph, and using its eigenvectors to define pseudo-reward functions for each Eigenoption to maximise. In Office, we used the first $32$ eigenvectors (and their negations). In all other environments, we used the first $16$ eigenvectors (and their negations).
Finally, we produced options for navigating to local maxima of betweenness Şimşek and Barto (2009). We computed the betweenness centrality of each node on the state-transition graph and selected all local maxima as subgoals. We defined options for navigating to each subgoal from the nearest $30$ states.
We arranged the Louvain options into a multi-level hierarchy, where options for navigating between clusters in partition $i$ were able to call skills for navigating between clusters in partition $i-1$, and only options at the lowest level of the hierarchy were able to call primitive actions. All of the options produced using alternative methods called primitive actions directly.
How we trained option policies. For all methods except Eigenoptions, we trained option policies using macro-Q learning McGovern et al. (1997). We used learning rate $\alpha=0.6$, initial action-values $Q_{0}=0$, and discount factor $\gamma=1$. All agents used an $\epsilon$-greedy exploration strategy with $\epsilon=0.2$.
For the clustering-based methods, we started the agent in a random state in the source cluster, and gave them rewards of $-0.01$ for each action taken and an additional $+1.0$ for reaching a state in the goal cluster. For node betweenness, we started the agent in a random state in the initiation set, and gave it a reward of $-0.01$ each decision stage, an additional $+1.0$ for reaching the subgoal state, and an additional $-1.0$ for leaving the initiation set. For Eigenoptions, we used value iteration to produce policies that maximised each Eigenoption’s pseudo-reward function.
How we trained our agents. We trained all agents using macro-Q and intra-option learning updates, which were performed every time an option at any level of the hierarchy terminated. We used a learning rate of $\alpha=0.4$, discount factor of $\gamma=1$, and initial action-values of $Q_{0}=0$ in all experiments. All agents used an $\epsilon$-greedy exploration strategy with $\epsilon=0.1$.
All learning curves we present show evaluation performance. After training each agent for one epoch, we evaluated the learned greedy policy by disabling exploration and learning (i.e., by setting $\epsilon=0$ and $\alpha=0$) and recording the rewards collected for one epoch in a separate instance of the environment.
To account for the stochasticity in the training process and the graph clustering algorithms used, we report the mean performance and standard error over $40$ runs.
How we trained the incremental agents. The training procedure for the incremental agents was largely identical to that of the non-incremental agents. The only difference is that the agent started with access to only primitive actions and, after a set number of decision stages, updated its state-transition graph and option hierarchy. A high-level overview of this training procedure is shown in Algorithm 2.
How we derived Louvain skills in Pinball. We used the existing graph construction method of Mahadevan and Maggioni (2007) to derive Louvain skills in the Pinball environment Konidaris and Barto (2009). We started by randomly sampling $4000$ states, and added edges between nodes representing each state and its $10$ nearest neighbours. We then assigned each edge $(u,v)$ a weight of $e^{-\nicefrac{{\lVert u-v\rVert^{2}}}{{\sigma}}}$ , with $\sigma=0.25$. Finally, we applied the Louvain algorithm ($\rho=0.05$) to the resulting graph.
Appendix G Comparison to Betweenness Skills
Here, we explore the similarities and differences between Louvain skills and skills that take the agent to local maxima of betweenness Şimşek and Barto (2009). This is a state-of-the-art subgoal-based approach that captures the bottleneck concept. Like us, this earlier work put forward an explicit skill characterisation, but presented only a single layer of skills defined over the primitive actions.
Many Louvain skills involve traversing bottlenecks in the state-space, such as navigating between rooms in a gridworld, or picking up the passenger and navigating between the walled sections in the Taxi domain. These bottlenecks are local maxima of betweenness. In Towers of Hanoi, all Louvain skills traverse states also identified as local maxima of betweenness: the highest local maxima of betweenness correspond to states separating the level 3 clusters, and the second-highest local maxima of betweenness correspond to states separating the level 2 clusters. However, not all Louvain skills traverse bottlenecks. In Rooms, lower levels of the cluster hierarchy divide each of the rooms into smaller regions, and skills for navigating between these regions do not correspond to traversing bottlenecks.
Louvain skills and betweenness skills also differ in how they can be arranged hierarchically. Even in Towers of Hanoi, where there is clearly hierarchical structure present between the larger and smaller local maxima of betweenness, it is unclear how to exploit the betweenness metric to form a multi-level hierarchy. In contrast, the modularity metric approximated by the Louvain algorithm provides us with a clear and principled way of building a multi-level hierarchy.
Appendix H Compute resource usage
Our experiments were run using a shared internal CPU cluster with specifications shown in Table 1.
Table 1: CPU cluster specifications
Processor 2$\times$ AMD EPYC 7443
Cores per Processor 24 Cores
Clock Speed 2.85GHz–4GHz
RAM 512 GB
RAM Speed 3200MHz DDR4
We utilised approximately 40 CPU cores for approximately 336 hours when producing the final set of results presented in the paper. Prior to this, we utilised approximately 20 CPU cores for approximately 168 hours during preliminary testing.
We did not make use of GPU acceleration because our experiments involved tabular reinforcement learning methods. We do not have access to estimates of our $\text{CO}_{2}$ emissions at this time.
0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 6 Discussion and Future Work
Joshua B. Evans
Department of Computer Science
University of Bath
Bath, United Kingdom
jbe25@bath.ac.uk
&Özgür Şimşek
Department of Computer Science
University of Bath
Bath, United Kingdom
o.simsek@bath.ac.uk
Appendix A Environments
Gridworlds include Rooms, Grid, Office, and Maze (Ramesh et al., 2019). They had four primitive actions: north, south, east, and west. These actions move the agent in the intended direction (unless the intended direction faces a wall, in which case the agent remains in the same state). The reward is $-0.001$ for each action and an additional $+1.0$ for reaching the goal state. There is a single start state and a single goal state, selected for each run from a list of possibilities.
Multi-Floor Office is an extension of Office to multiple floors. All foors are connected by an elevator, which occupies the same grid square on each floor, where the agent has two additional primitive actions: up and down. These actions move the agent to the adjecent floor in the intended direction (unless there is no other floor in that direction).
Taxi is a $5\times 5$ grid with four special squares (R, G, B, and Y) that serve as possible pick-up and drop-off locations for a passenger. An episode starts with the taxi at a random square, the passenger at a random special square, and the destination another random special square. Six actions are available in each state: north, south, east, west, pick-up, and put-down. The navigation actions are identical to those in the gridworlds, as described above. Pick-up and put-down have the intended effect when appropriate; otherwise they do not change the state. The reward is $-0.001$ for each action and an additional $+1.0$ when the passenger is put down at the destination.
Towers of Hanoi contains four discs of different sizes, placed on three poles. An episode starts with all discs stacked on the leftmost pole. Primitive actions move the top disc from one pole to any other pole, with the constraint that a disc cannot be placed on top of a disc smaller than itself. The reward is $-0.001$ for each action and an additional reward of $+1.0$ at the goal state, where when all three discs are stacked on the rightmost pole.
Appendix B Methodology
Generating options. To generate Louvain options, the Louvain algorithm ($\rho=0.05$) was applied to the state transition graph, resulting in a set of partitions. Any partition where the mean number of nodes per cluster was smaller than $4$ was discarded. For all remaining partitions, options were defined for efficiently taking the agent from a cluster $c_{i}$ to each of its neighbouring clusters $c_{j}$ if a directed edge existed from a node in $c_{i}$ to a node in $c_{j}$. The Louvain options were arranged into a multi-level hierarchy, where options for navigating between clusters in partition $i$ could call skills for navigating between clusters in partition $i-1$. Only options at the lowest level of the hierarchy could call primitive actions. Options generated using alternative methods called primitive actions directly.
Eigenoptions (Machado et al., 2017) were generated by computing the normalised Laplacian of the state transition graph and using its eigenvectors to define pseudo-reward functions for each Eigenoption to maximise. In Office, the first $32$ eigenvectors (and their negations) were used. In all other environments, the first $16$ eigenvectors (and their negations) were used. Options for navigating to local maxima of betweenness Şimşek and Barto (2009) were generated by selecting all local maxima as subgoals and defining options for navigating to each subgoal from the nearest $30$ states.
Learning option policies. For all methods except Eigenoptions, option policies were learned using macro-Q learning McGovern et al. (1997), with learning rate $\alpha=0.6$, initial action values $Q_{0}=0$, and discount factor $\gamma=1$. All agents used an $\epsilon$-greedy exploration strategy with $\epsilon=0.2$. For options based on clustering, the agent started in a random state in the source cluster. It received a reward of $-0.01$ for each action taken and an additional $+1.0$ for reaching a state in the goal cluster. For options based on node betweenness, the agent started in a random state in the initiation set. It received a reward of $-0.01$ at each decision stage, an additional $+1.0$ for reaching the subgoal state, and an additional $-1.0$ for leaving the initiation set. For Eigenoptions, value iteration was used to produce policies that maximised the pseudo-reward function of each Eigenoption.
Producing learning curves. All agents were trained using macro-Q McGovern et al. (1997) and intra-option learning Sutton and Precup (1998) updates, which were performed every time an option at any level of the hierarchy terminated. A learning rate of $\alpha=0.4$, discount factor of $\gamma=1$, and initial action values of $Q_{0}=0$ were used in all experiments. All agents used an $\epsilon$-greedy exploration strategy with $\epsilon=0.1$. All learning curves show evaluation performance. After training each agent for one epoch, the learned policy was evaluated (with exploration and learning disabled) in a separate instance of the environment.
Applying the Louvain algorithm in Pinball. The state transition graph was constructed by following the method used by Mahadevan and Maggioni (2007). Initially, $4000$ states were randomly sampled and a node representing each state was added to the graph. Edges were then added between each node and its $10$ nearest neighbours. An between two locations $u$ and $v$ with Euclidean distance $d$ between them was assigned a weight of $e^{-4d^{2}}$. Finally, the Louvain algorithm ($\rho=0.05$) was applied to the resulting graph.
Appendix C Comparison to Skills that Navigate to Local Maxima of Betweenness
Here we explore the relationship between Louvain skills and the skill chracterisation proposed by Şimşek and Barto (2009). The latter is a subgoal-based approach that captures various definitions of the bottleneck concept. It defines skills that navigate to local maxima of betweenness. Both approaches address the conceptual problem of what makes a useful skill, explicitly defining a target set of skills for the agent to learn.
There is substantial overlap between Louvain skills and skills that navigate to local maxima of betweenness. They both include skills that traverse bottleneck states, including those that navigate between rooms in Rooms, picking up the passenger in Taxi, and navigating different parts of the grid in Taxi. In Towers of Hanoi, all Louvain skills traverse states that are also identified as local maxima of betweenness. The highest local maxima of betweenness correspond to the states that separate Louvain clusters at level 3; the second highest local maxima of betweenness correspond to the states that separate Louvain clusters at level 2.
On the other hand, there are many Louvain skills that do not correspond to navgating to local maxima of betweenness. Examples include the Louvain skills that navigate within a single room in Rooms.
Most importantly, Louvain skills and skills that navigate to local maxima of betweenness differ in how they can be arranged hierarchically. Even in Towers of Hanoi, where there is a clear hierarchical structure between the larger and smaller local maxima of betweenness, it is not clear how to exploit the betweenness metric to form a multi-level hierarchy. In contrast, the modularity metric approximated by the Louvain algorithm provides a clear and principled way of building a multi-level hierarchy.
Appendix D Cluster Hierarchies
Figure 1 shows the cluster hierarchies produced by the Louvain algorithm when applied to the state transition graphs of Grid and Maze. It also shows the first level of the cluster hierarchy in Office, which was omitted in the main paper due to space limitations.
Level 4
Level 3
Level 2
Level 1
Level 5
Level 4
Level 3
Level 2
Level 1
Level 1
Figure 1: Top two rows: Cluster hierarchies produced by the Louvain algorithm in Grid and Maze. Bottom row: The lowest level of the cluster hierarchy in Office.
Appendix E Sensitivity to the Resolution Parameter
@@
Figure 2 shows the performance of agents in Rooms and Towers of Hanoi using Louvain skills generated using different values of $\rho$. Louvain skills created using lower values of $\rho$ consistently outperformed those created by using higher values, and very high values ($\rho\geq 10$) generally led to performance similar to that obtained by using primitive actions only. Lower $\rho$ values lead to deeper hierarchies that contain skills for navigating the environment over varying timescales. In contrast, higher $\rho$ values result in shallower skill hierarchies that contain few—or, in the extreme, no—levels of skills above primitive actions. While there are better and worse values of $\rho$, it may be argued that there is no “bad” choice; lowering $\rho$ will result in deeper hierarchies, but existing levels of the hierarchy will remain intact.
Figure 2: Agent performance with Louvain skills generated using different settings of the resolution parameter $\rho$.
Appendix F Compute Resource Usage
The experiments were run using a shared internal CPU cluster with the specifications shown below. Approximately 40 CPU cores were utilised for approximately 336 hours when producing the final set of results presented in the paper. Prior to this, approximately 20 CPU cores were utilised for approximately 168 hours during preliminary testing. GPU acceleration was not used because the experiments involved tabular reinforcement learning methods.
Processor 2$\times$ AMD EPYC 7443
Cores per Processor 24 Cores
Clock Speed 2.85GHz–4GHz
RAM 512 GB
RAM Speed 3200MHz DDR4
Appendix G Source Code
An implementation of the proposed approach, implementations of the test environments, and the code used to generate all results presented in the paper can be found in the following GitHub repository: https://github.com/bath-reinforcement-learning-lab/Louvain-Skills-NeurIPS-2023
Appendix H Louvain Algorithm
@@
Algorithm 1 shows pseudocode for the Louvain algorithm (Blondel et al., 2008). It takes a graph $G_{0}=(V_{0},E_{0})$ as input and outputs a set of partitions of that graph into clusters. Each iteration of the algorithm (lines 4–25) produces one partition of the graph.
@@
1 parameters: resolution parameter $\rho\in\mathbb{R}^{+}$ input: $G_{0}=(V_{0},E_{0})$ // e.g., the state transition graph of an MDP $i\leftarrow 0$ repeat
@@ 8             $i\leftarrow i+1$
Algorithm 1 Louvain Algorithm
Appendix I Incremental Discovery of Louvain Skills
Algorithm 2 presents an approach for incrementally developing a hierarchy of Louvain skills over time, starting with only primitive actions. To update the agent’s skill hierarchy, Algorithm 2 calls upon either Algorithm 1 or Algorithm 3, depending on whether the agent’s existing skill hierarchy is to be replaced or updated. Algorithm 3 presents an approach for incrementally updating Louvain partitions. It integrates new nodes into an existing cluster hierarchy.
@@ Appendix I Incremental Discovery of Louvain Skills
$G\leftarrow(V,E)$   // initialise empty state transition graph (STG)
@@ 20            $t\leftarrow t+1$
Algorithm 2 Incremental Discovery of Louvain Skills
parameters: resolution parameter $\rho\in\mathbb{R}^{+}$ input: $G_{0}=(V_{0},E_{0})$   // e.g., the state transition graph (STG) of an MDP
@@ 8             $i\leftarrow i+1$
Algorithm 3 Update Louvain Partitions